A lawnmower takes 750 complete revolutions to cut grass on a field. Calculate the area of the field if the diameter of the lawnmower is 84 cm and length is 1 m.

 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

length of lawnmower = 1m = 100cm

Its circumference = π × D = 22/7 × 84 = 264 cm

Length of field will be = 264 × 750 = 198000 cm

Here, the width of field = length of the lawnmower i.e. 100 cm

So, area of field = 198000 × 100 = 19,800,000 cm²

= 1980 m²

Thanks
Daya kaur
one year ago
The area of a square ground is 64084 sq.m. it is fenced at the rate of ₹40 per m and ploughed at the rate of ₹20per sq.m find the total coast